1. How many moles of NaCl are contained in 100.0 mL of a 0.20 M solution?  What mass is this?

 




1
Expert's answer
2021-05-06T07:39:14-0400

The number of moles can be calculated as follows:

n = M × V

where n - the number of moles, M - molarity, V - volume.

From here:

n = 0.20 M × 100.0 mL = 0.20 M × 0.100 L = 0.020 mol

The mass of the solute equals:

m = n × Mr(NaCl) = 0.020 mol × 58.44 g/mol = 1.17 g


Answer: 0.020 mol; 1.17 g.